OpenStudy (broskishelleh):

Two terms: y = –2x^2 + 6 y = -0.1x^2 + 6

OpenStudy (broskishelleh):

You are basically making a coaster Create an equation that will be steeper than both equations, and explain your reasoning.Is what I needed help with

OpenStudy (broskishelleh):

I know they are both conics y = (-1/10) x2 + 6 has the lines of the conic touch -8 and 8 on the x-axis, with the conic from question A has it touching -2 and 2 on the x - axis

OpenStudy (broskishelleh):

@aum @phi @mathmale @mathstudent55

OpenStudy (broskishelleh):

@ikram002p

OpenStudy (dbzfan836):

@Destinymasha

OpenStudy (broskishelleh):

@SolomonZelman

OpenStudy (broskishelleh):

@ganeshie8 @phi

OpenStudy (broskishelleh):

@undeadknight26

OpenStudy (broskishelleh):

@PFEH.1999

OpenStudy (broskishelleh):

Thank you for coming

OpenStudy (phi):

"steeper" means a small step in x causes a large step in y y = –2x^2 + 6 here, if we move from x=0 (y=6) to x=1 (y=4) y goes down by 2 if we put a bigger number in front of the x, it gets steeper y = -4x^2 + 6 at x=0 y = 6 at x=1 y = 2 we go down by 4, which is a steeper drop

OpenStudy (broskishelleh):

This is the answer?

OpenStudy (phi):

Create an equation that will be steeper than both equations, that means make the coefficient of the x^2 term bigger (in absolute terms)

OpenStudy (broskishelleh):

Alright then So like you said y = -4x^2 + 6 for the first one y = -(1/20)x^2 + 6

OpenStudy (aum):

I think they just want ONE equation that is steeper than BOTH.

OpenStudy (broskishelleh):

So y = -4x^2 + 6

OpenStudy (aum):

That will do. But you have to explain your reasoning. In y = ax^2 +b, increasing the magnitude of a, that is the absolute value of a, has the effect of stretching the graph vertically, making it steeper.

OpenStudy (broskishelleh):

Thanks @aum

OpenStudy (aum):

You are welcome.